OpenStudy (usernamenottaken):

Simplify (5- 3i)^2. Will medal.

OpenStudy (sweetburger):

Expand it first using foil method. writing it like this (5-3i)(5-3i)

jhonyy9 (jhonyy9):

exactly @sweetburger nice job

OpenStudy (usernamenottaken):

so 16 - 30 i

OpenStudy (usernamenottaken):

anyone?

jhonyy9 (jhonyy9):

(5-3i)(5-3i) = 5*5 -5*3i -3i*5 -3i*(-3i) = ?

jhonyy9 (jhonyy9):

step by step not directly answers - for anybody can understanding it right sure easy

OpenStudy (usernamenottaken):

thats what i got, i even double checked with wolframalpha http://www.wolframalpha.com/input/?i=(5-3i)(5-3i)

jhonyy9 (jhonyy9):

(5-3i)(5-3i) = 5*5 -5*3i -3i*5 -3i*(-3i) = ? =25 -15i -15i +9(i^2) = =25 -30i + ?

jhonyy9 (jhonyy9):

do you can continue it please ?